While taking a trip to Seattle, a plane flies against the wind 288 miles and then returns home with the same wind. In calm wind, the plane can fly 300mph. If the total flying time was 2 hours, what is the speed of the wind?
Answer by mananth(16946)

While taking a trip to Seattle, a plane flies against the wind 288 miles and then returns home with the same wind. In calm wind, the plane can fly 300mph. If the total flying time was 2 hours, what is the speed of the wind?
let wind speed be x mph
headwind speed = 300-x
tailwind speed = 300+x
distance =288
Time =2 hours
time to seattle and return =2
288/(300-x) + 288/(300+x)=2
(288(300-x) +288(300+x))/((300+x)(300-x))=2
(288(300-x) +288(300+x)) = (90000-x^2)*2


86400-288x +86400 +288x= 180000-2x^2
2x^2= 7200
x^2 =3600
x= 60 mph wind speed